OpenStudy (anonymous):

Find the volume of a cylinder with a height of 2 meters and a diameter that is 5 times the measure of the height. Use 3.14 for pi, and round your answer to the nearest hundredth 157 339.12 678.24 1570

OpenStudy (anonymous):

V=πr^2h

OpenStudy (anonymous):

V = 3.14 * 5 * 2

OpenStudy (anonymous):

They try to trick you by saying 'D is 5 times the measure of the height'

zepdrix (zepdrix):

Just for clarity:\[\Large\rm h=2\]\[\Large\rm d=5h=10\]\[\Large\rm r=\frac{1}{2}d=5\]That's where jc got those numbers from.

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

No problem.